### v2

# InSTA: Towards Internet-Scale Training For Agents

Abstract
The predominant approach for training web navigation agents is to gather human demonstrations for a set of popular websites and hand-written tasks, but it is becoming clear that human data is an inefficient resource. We develop a pipeline to facilitate internet-scale training for agents without laborious human annotations. In the first stage, an LLM annotates 150k sites with agentic tasks. In the next stage, LLM agents complete tasks and produce trajectories. In the final stage, an LLM filters trajectories by judging their success. Language models are powerful data curation tools, identifying harmful content with an accuracy of 97%, judging successful trajectories with an accuracy of 82.6%, and producing effective data. We train agents based on Qwen 3 1.7B that are competitive with frontier LLMs as web agents, while being smaller and faster. Our top agent reaches a success rate of 56.9%, outperforming the data collection policy Qwen 3 235B, a 235 times larger Llama 4 Maverick, and reaching 94.7% of the performance of Gemini 2.5 Flash. We are releasing code, models and data at: https://data-for-agents.github.io.
1 Introduction
The predominant approach for training LLM web navigation agents is to collect human demonstrations for a set of manually curated websites and tasks (Deng et al., 2023; Zhou et al., 2024b; Putta et al., 2024; Koh et al., 2024a; Liu et al., 2024; Lù et al., 2024; Rawles et al., 2023). Human data can be laborious to collect, and becomes costly to scale as the breadth of skills that users require from language model agents grows. There are more than 300M sites on the western internet according to The Common Crawl Foundation (2025), and the range of sites that researchers have annotated represents a tiny fraction of the vast available data. And crucially, the existing human data is static. There is a growing need to automate pipelines for training the next generation of language model agents in a dynamic internet-scale environment. This paper addresses the core challenge of building this environment—reducing dependence on human annotations. We develop an automatic pipeline that aims to facilitate internet-scale training for agents, which we refer to as the InSTA pipeline.
Figure 1: Overview of the InSTA pipeline. Our work unlocks a dynamic internet-scale environment that allows training small models to match frontier LLMs as agents, on a fraction of the budget. Starting from the top 1M sites on the internet, we annotate 150k sites with challenging agentic tasks, and release the entire pipeline, including code, models and an official huggingface dataset, on our website: https://data-for-agents.github.io.
The InSTA pipeline has three stages. In the first stage, we employ a language model task proposer that annotates 150k sites with live web navigation tasks for agents to perform. Existing works are limited to 200 popular websites (He et al., 2024; Lù et al., 2024; Rawles et al., 2023; Deng et al., 2023; Zhou et al., 2024c; Murty et al., 2025) that researchers have annotated manually, and a handful of synthetic websites (Zhou et al., 2024b; Koh et al., 2024a; Yao et al., 2023a). Our first goal in this paper is to improve coverage of real-world sites. To accomplish this, we cast a wide net. Starting from the top 1M sites on the internet ranked by popularity, our task proposer filters down to 150k sites that have safe content. Safety is critical when building autonomous agents, and our task proposer succeeds at detecting harmful content with an accuracy of 97%. Tasks are generated for sites marked as safe by the task proposer, and we run language model agents to complete the generated tasks. Agent progress is then fed back to the task proposer, which reviews trajectories and judge evaluations in order to assign a harder task based on the latest content of the website, closing the loop.
Scaling the task generation loop, we annotate 150k diverse sites with challenging agentic tasks, and release an official huggingface dataset: data-for-agents/insta-150k-v2. Motivated by the importance of internet data to progress in modern deep learning, our second goal in this paper is an internet-scale data flywheel for training LLM agents. We approach this by harnessing LLMs as data curation tools. After generating tasks, the pipeline employs pretrained LLMs as agents to complete tasks and produce trajectories, which are filtered by a judge to select the best data. Agents control a virtual web browser and produce reasoning traces that contain function calls to interact with and navigate live webpages. The judge produces a reasoning trace that considers whether a trajectory is successful, and scores the agent on a continuous scale from 0 to 1. We scale the pipeline to create a large reasoning dataset for multimodal agents, with 2.2M screenshots, 2.2M traces for actions, and 150k traces for the judge.
Our data unlocks great potential in small language models as agents. We train a series of models based on Qwen 3 1.7B on varying scales of data from the InSTA pipeline, and match the performance of frontier LLM agents, on a fraction of the budget. Our top agent has a success rate of 56.9%, outperforming the data collection policy Qwen 3 235B, beating a 235 times larger Llama 4 Maverick, and reaching 94.7% of the performance of Gemini 2.5 Flash. To share our progress, we are releasing the entire pipeline, including code, models and data, on our website: https://data-for-agents.github.io.
2 Related Works
Language Model Agents.
There is an emerging paradigm in modern NLP using language models (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a, b) as the backbone for agents (Andreas, 2022). These models show impressive reasoning capabilities (Bubeck et al., 2023; Zhong et al., 2024; Valmeekam et al., 2024) that allow them to generalize to downstream applications, such as web navigation, where observations differ from LLM training data. Search algorithms provide a secondary axis to improve the reasoning capabilities of language model agents (Yao et al., 2023b; Besta et al., 2024; Koh et al., 2024b; Zhou et al., 2024a) by providing an explicit algorithmic scaffold and allowing test-time compute to improve reasoning steps (Snell et al., 2024; Zhong et al., 2024). Although most recent work focuses on running language models as zero-shot agents, fine-tuning language models to improve their effectiveness as agents is becoming popular (Putta et al., 2024; Zeng et al., 2023; Zhang et al., 2023; Hong et al., 2023; Xie et al., 2024; Wang et al., 2024; Zhou et al., 2024c; Murty et al., 2025) as target benchmarks are becoming more difficult to solve zero-shot.
Agent Pipelines.
There are a growing number of agent pipelines aimed at fine-tuning language models to improve their effectiveness as agents (Mitra et al., 2024; Zeng et al., 2023; Putta et al., 2024; Chen et al., 2023; Ou et al., 2024; Zhou et al., 2024c; Murty et al., 2025). However, driven by the limited data available, many such works train on data with significant overlap with their test environment, either with different tasks for the same environment configuration as the test setting (Deng et al., 2023; Zhou et al., 2024c; Murty et al., 2025), or even the same tasks (Putta et al., 2024). We consider a setting where tasks and environment configurations (i.e. websites) are entirely separate between training and testing, creating a strong train-test split. This presents a challenge: human data for training LLM agents is limited (Deng et al., 2023; Lù et al., 2024). We address this challenge by reducing dependence on human data in agent pipelines. We employ LLMs as data curation tools to automatically design challenging tasks, and select the best training data. Our pipeline allows us to train small models that match frontier LLMs on Web Voyager (He et al., 2024), without using any data from Web Voyager. Contrast this with methods that train primarily on data from Web Voyager (Zhou et al., 2024c; Murty et al., 2025), and may not transfer to other benchmarks (Xue et al., 2025).
Agent Datasets.
Datasets for training web navigation agents typically rely on human annotators to create tasks (Zhou et al., 2024b; Koh et al., 2024a; Rawles et al., 2023), and provide demonstrations for tasks (Deng et al., 2023; Lù et al., 2024; Rawles et al., 2023; Shen et al., 2024). However, the amount of data researchers have annotated represents a tiny fraction of the available internet data. There are more than 300M sites on the internet according to The Common Crawl Foundation (2025), yet existing datasets are limited to 200 popular sites that human annotators are familiar with (Deng et al., 2023; Lù et al., 2024; Shen et al., 2024). Human data can be laborious to collect, and becomes costly to scale as the capabilities users require from agents grows. And crucially, human data is static. Our work moves away from fixed datasets for training agents, and towards a dynamic internet-scale environment that grows with an ever-changing internet. We are not the first to build an environment (Zhou et al., 2024b; Koh et al., 2024a; Yao et al., 2023a; He et al., 2024), nor are we the first to consider synthetic data (Gandhi et al., 2024; Ou et al., 2024; Setlur et al., 2024; Tajwar et al., 2024), but we have solved a key challenge that unlocks the internet as the largest environment for agents.
Language Model Judges.
Using LLMs to judge the correctness of responses is becoming popular to refine LLM predictions (Li et al., 2024), including to verify reasoning steps (Zhang et al., 2024), rejection sample (Snell et al., 2024; Sun et al., 2024), prioritize frontier nodes in search algorithms (Zhou et al., 2024a; Koh et al., 2024b), filter out harmful responses (Inan et al., 2023), provide feedback for response improvement (Madaan et al., 2023; Paul et al., 2024; Patel et al., 2024; Yuksekgonul et al., 2024), and provide ratings for alignment (Lee et al., 2024; Ouyang et al., 2024). Our use of language models to evaluate agents is inspired by Generative Verifiers (Zhang et al., 2024), and the multimodal verifier in He et al. (2024). One difference is our verifier predicts a reasoning trace that scores the agent from 0 to 1, which helps us rank trajectories to select the best data.
3 Language Model Agents
Language model agents are a class of decision-making agents represented by $\pi_{\text{LLM}}(\mathbf{a}_{t}|\mathbf{s}_{t},\mathbf{c})$, a policy that processes multimodal observations $\mathbf{s}_{t}$ (from a virtual web browser in our case) and predicts textual actions $\mathbf{a}_{t}$ to complete a task $\mathbf{c}$. Underneath this abstraction, a large language model (LLM) generates actions via the next-token prediction, conditioned on a system prompt $\mathbf{x}_{\text{agent}}$.
$\mathbf{a}_{t}=f^{\;\text{text}\to\text{act}}(\;\text{LLM}(\;\mathbf{x}_{\text{agent}},\mathbf{c},\text{Enc}(\mathbf{s}_{t})\;)\;)$ (1)
Environment representations for observations and actions typically differ from the expected input format of the language model (typically images and text), and functions are introduced that map the observations to a multimodal prompt $\text{Enc}(\cdot)$, and parse actions from the language model generated response $f^{\;\text{text}\to\text{act}}(\cdot)$. For web navigation, the environment state $\mathbf{s}_{t}$ is HTML DOM, and is often formatted as raw HTML code, an Accessibility Tree, Set-of-marks, or screenshots (Zhou et al., 2024b; Koh et al., 2024a; Chezelles et al., 2024; Shen et al., 2024). We built a fast Markdown parser that converts webpage observations into a compact readable format (refer to the code). Action formats vary between works, and we build on Schick et al. (2023)’s function-calling framework, where a language model generates code that is parsed into a function name and corresponding arguments. Given sets of function and argument names $L_{i}$, and sets of argument values $G_{i}$, the action space $\mathcal{A}$ is:
$\mathcal{A}=L_{\text{func}}\times(L_{\text{arg1}}\times G_{\text{arg1}})\times(L_{\text{arg2}}\times G_{\text{arg2}})\times\cdots\times(L_{\text{argN}}\times G_{\text{argN}})$ (2)
where $L_{\text{func}}$ is the set of function names on the page object in the Playwright API (Microsoft, 2024), and function arguments have a name and value ($L_{\text{arg1}}\times G_{\text{arg1}}$) corresponding to the Playwright API. We allow the agent access to call arbitrary functions in Playwright (Microsoft, 2024), a Microsoft-developed browser automation library that wraps a headless web browser. The agent’s goal is to complete a web navigation task specified via a natural language instruction $\mathbf{c}\in L$, starting from an initial URL, and operating the browser via function calls to the Playwright API until the desired task is complete, after which the agent calls the stop function and provides a final response:
[…]
We prompt the agent to produce a reasoning trace of a desired length (ablated in Figure 10) that contains function call actions as JSON in a fenced code block. To parse actions from the response, we employ a regex template that matches the first JSON code block, and a JSON decoder $f^{\;\text{text}\to\text{act}}(\cdot)$ to parse the contents within the code block. When parsing fails due to invalid syntax, we allow the agent to generate a second response. Equipped with a language model agent that makes function calls with the Playwright API, we may consider the crucial task of obtaining large and diverse data.
4 Internet-Scale Task Generation
Training the next generation of LLM agents requires a large and diverse set of websites and tasks beyond what researchers have annotated so far (Deng et al., 2023; Zhou et al., 2024b; Koh et al., 2024a; Liu et al., 2024; Lù et al., 2024; Rawles et al., 2023; He et al., 2024). We develop an approach to efficiently annotate vast numbers of sites from diverse sections of the internet with agentic tasks. Our approach introduces two important desiderata: (1) it should not rely on human annotations, and (2) tasks should derive from a feedback process that deeply explores the environment.
4.1 Language Model Task Proposer
The key idea in stage one of the pipeline is a feedback loop, where a language model task proposer $\psi_{\text{LLM}}(\cdot)$ guides exploration on a website via an initial easy task. We then run a language model agent to explore the site, conditioned on the initial task, which produces an exploratory trajectory that is fed back to the task proposer. Conditioned on a trajectory that deeply explores the website, the task proposer then creates a harder, grounded task. This process is summarized as an equation.
$\mathbf{c}\sim\psi_{\text{LLM}}(\mathbf{c}|\mathbf{w},\tilde{\mathbf{c}},\underbrace{\mathbf{s}_{1},\mathbf{a}_{1},\cdots,\mathbf{s}_{T},\mathbf{a}_{T}}_{\text{trajectory}})\;\;\text{st}\;\;\mathbf{a}_{t}\sim\pi_{\text{LLM}}(\mathbf{a}_{t}|\mathbf{s}_{t},\tilde{\mathbf{c}})\;\;\mathbf{s}_{t+1}\sim\mathbb{P}(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t})$ (4)
In this equation, the website url is $\mathbf{w}$, the initial task is $\tilde{\mathbf{c}}$, the trajectory includes the states $\mathbf{s}_{t}$ and the actions $\mathbf{a}_{t}$, and the harder, grounded task is $\mathbf{c}$. Highlighted in Figure 2, we annotate 150k sites with tasks by scaling Equation 4, and release them on huggingface at: data-for-agents/insta-150k-v2.
Figure 2: Annotating 150k live sites with agentic tasks. Starting from 1,000,000 websites, we employ a pretrained language model that marks sites as safe/unsafe for annotation, and assigns a realistic task that a hypothetical user might want to accomplish on each site. The task proposer aggressively filters out 85% of websites from the pipeline, resulting in 150k safe websites annotated with realistic tasks.
Model Details.
We utilize pretrained and frozen language models that conform to a chat interface and accept system, user, and assistant prompts. The task proposer system prompt is listed in Appendix 11, and details all cases for which sites are considered unsafe. We employ the Llama 3.1 family of LLMs from Meta (Grattafiori et al., 2024; Touvron et al., 2023b, a), the GPT family of LLMs from OpenAI, and the Gemini family of LLMs from Google. Inference is served using vLLM (Kwon et al., 2023) for the Llama series of models. We employ a sampling temperature of 0.5 and a maximum budget of 1024 new generated tokens, while all other parameters are kept as defaults in the OpenAI chat completions API, which is used to make inference calls to all LLMs.
Prompt Details.
[…]
The design of the task proposer as a feedback process is important, and the full potential of this design will be realized in future work that trains agents with an on-policy reinforcement learning algorithm. In such future work, the task proposer can be used within the RL loop to generate incrementally harder tasks as agents learn. For this paper, we employ one loop of task generation.
Figure 3: Most frequent words in our tasks. This wordcloud shows the top 500 most frequent words in tasks from the training set of our official huggingface dataset. The size of each word corresponds to its frequency in the dataset. Our tasks span diverse categories and lexicon.
4.2 Safety & Reliability
Safety is critical when building autonomous agents. The internet contains significant amounts of content that should be removed from training data, in order to avoid agents learning harmful behaviors. To understand the robustness of our safety filter, we input 100 carefully selected websites to the task proposer, of which 50 contain harmful, or mature content. Table 2 reports the accuracy, precision, and recall of the safety filter on this data. For a variety of LLMs, the safety filter displays high accuracy—up to 97% of websites are correctly classified, and recall for detecting unsafe websites is as high as 1.0, suggesting that nearly all unsafe websites are detected by the safety filter.
Method Acc. Prec. Recall
Llama 3.1 70B 85% 0.77 1.00
GPT-4o 95% 0.91 1.00
Gemini 1.5 Pro 97% 0.96 0.98
Table 1: The safety filter displays a high accuracy. We measure the accuracy, precision, and recall of the safety filter on a set of 100 websites, where 50 contain harmful, or mature content. Up to 97% of websites are correctly classified, and recall is as high as 1.0.
Method Verifiable Rate
Llama 3.1 70B 75%
GPT-4o 85%
Gemini 1.5 Pro 89%
Table 2: Generated tasks are typically achievable. We measure the rate that human workers were able to complete and verify their completion of tasks produced by the task proposer for a set of 100 websites. Up to 89% of tasks are achievable, and verifiable.
Reliability is equally important for autonomous agents. Instructions should be followed faithfully by agents, and this requires training them with tasks that are achievable, and verifiable. Table 2 reports the rate that human workers were able to complete and verify their completion of tasks produced by the task proposer in its initial phase. Up to 89% of tasks are achievable, and verifiable according to the study, which suggests the pipeline is producing reliable tasks. Together with results in Section 6, it is likely that data from the InSTA pipeline leads to agents that follow instructions faithfully.
4.3 Scaling To 150,000 Websites
We leverage Common Crawl for task generation. As of May 2025, the latest web graph released by The Common Crawl Foundation (2025) contains more than 300M unique hosts, which we adapt into a data source for agents. In particular, we select the top 1M sites based on their PageRank values. Common Crawl likely contain many unsafe websites, and these are filtered out by the task proposer. Each phase of task generation requires 14 hours of compute time using two 8-GPU v100 machines, and filters to 150k safe websites annotated with tasks. Statistics of tasks are shown in Figure 3.
5 Internet-Scale Environment
By this point, we have reached our first goal—to improve coverage of real-world sites by annotating 150k diverse sites with challenging agentic tasks. To reach our second goal, and move beyond a static dataset, towards a dynamic internet-scale environment, we require a robust evaluator for these tasks. Evaluation presents a subtle challenge. The web evolves constantly, and daily changes in website content may invalidate a fixed ground truth reference solution. Driven by necessity, this environment must be evaluated by a model that judges whether an agent’s solution is correct, in the latest context.
Figure 4: Automatic evaluation for agents with language model judges. Building on the large and diverse set of tasks generated by the pipeline, we employ pretrained language models to attempt and evaluate web navigation tasks. We dispatch language model agents to perform tasks by making calls to the Playwright API. We then employ language model judges to evaluate the trajectories.
5.1 Evaluation With Language Models
We model the process of evaluating trajectories from agents as a classification problem, where the goal is to estimate the probability $\mathbf{r}_{T}$ that a task $\mathbf{c}$ is solved, and generate $\mathbf{r}_{T}$ via next-token prediction, conditioned on a system prompt $\mathbf{x}_{\text{judge}}$, a task $\mathbf{c}$, and a trajectory $\mathbf{s}_{1},\mathbf{a}_{1},\cdots,\mathbf{s}_{T},\mathbf{a}_{T}$. The LLM is instructed to produce a reasoning trace that scores the agent on a scale from 0 to 1, and embed the score as JSON in a fenced code block. We employ a regex template that matches to the first code block in the response, and a JSON decoder to parse $\mathbf{r}_{T}$ from the response, given by $f^{\;\text{text}\to\text{val}}(\cdot)$.
$\mathbf{r}_{T}=f^{\;\text{text}\to\text{val}}(\;\text{LLM}(\;\mathbf{x}_{\text{judge}},\mathbf{c},\text{Enc}(\mathbf{s}_{1}),\mathbf{a}_{1},\cdots,\text{Enc}(\mathbf{s}_{T}),\mathbf{a}_{T}\;)\;)$ (5)
Verifying The Judge.
To understand the robustness of the judge, we measure its accuracy detecting successful trajectories that were annotated by human workers. We annotate 100 trajectories with binary success labels, and apply a threshold $\mathbf{r}_{T}>0.5$ to obtain binary predictions from the judge. Figure 5 reports the accuracy of the judge as a function of the PageRank values of sites, and as a function of the confidence of the judge, given by conf = $2\cdot|\mathbf{r}_{T}-1/2|$. For all LLMs tested, the judge shows a high accuracy, ranging from $78.0\%$ for Gemini 1.5 Pro, to $81.7\%$ for Llama 3.1 70B, and $82.6\%$ for GPT-4o. Accuracy is stable as PageRank falls, suggesting the judge is accurate for less popular sites that may not be well represented in the LLM’s training data. Shockingly, the confidence predicted by LLMs is highly interpretable, and correlates with accuracy to the point that trajectories where conf = 1 are classified with an accuracy up to $93.1\%$. The emergent robustness of the judge equips us to efficiently and accurately verify agent solutions on a dynamic internet.
Figure 5: Language models are robust evaluators. We measure the accuracy of language models for detecting successful trajectories, and find that accuracy remains stable relative to PageRank values (left plot). As models become more confident, their accuracy improves (right plot), suggesting confidence is a useful proxy for the reliability of their predictions.
Figure 6: Statistics for our agent reasoning dataset. We conduct a large data collection experiment using our top checkpoint for Qwen3 1.7B. Our dataset has 2.2M screenshots, 2.2M reasoning traces for actions, and 150k traces for the judge. 50.0% of the trajectories are successful according to the judge, and have diverse ratings for efficiency and self-correction.
5.2 Scaling To 150,000 Agents
With our task proposer, agent, and judge driven by pretrained language models, we have all components needed to harness internet-scale data. We conduct a large data collection experiment, running language model agents to complete tasks on 150k websites from our official dataset, producing 2.2M screenshots and 2.2M reasoning traces for actions within 150k trajectories. The judge annotates these trajectories, producing 150k reasoning traces for evaluations, and leading to the statistics in Figure 6. For this experiment, we employ a fine-tuned Qwen3 1.7B as the agent (refer to the next section), and Qwen3 235B zero-shot as the judge. Data collection requires 1,200 v100 GPU hours, and costs \$521.55 based on current AWS spot instance pricing, a fraction of the budget that industry labs are spending towards agents. If you have the right data, this small budget is sufficient and no human annotations are required to build models that compete with frontier LLMs as agents.
6 Training Agents
Figure 7: InSTA unlocks great potential in small models. We train agents based on Qwen 3 1.7B using trajectories produced by a Qwen 3 235B data collection policy, and optionally filtered by a Qwen 3 235B judge (see Judge Filtered vs. Uniformly Sampled). We report success rates on a test set of 3,000 held-out websites and tasks. Before training, Qwen 3 1.7B has a zero-shot success rate of 11.5% according to a Qwen 3 235B judge, and we improve this by +45.3% absolute percentage points. Our top checkpoint outperforms the Qwen 3 235B data collection policy, and Llama 4 Maverick, a frontier LLM with 400B parameters, for which our model is 235 times smaller. Notably, filtering with a Qwen 3 235B judge leads to agents that improve according to independent secondary judges, including Gemini 2.5 Flash, Llama 4 Maverick, and Gpt 4.1 Nano, suggesting it generalizes well.
We’ve built an internet-scale data flywheel that produces trajectories annotated with scores from a judge that can help us train LLM agents. To understand the quality of data produced by the flywheel, we conduct a series of experiments training models on the data, and testing on popular benchmarks. These experiments focus on three main questions: (1) what is the impact of increasing data scale? (2) do agents transfer to new domains? (3) do agents scale with test-time compute?
6.1 Performance Improves With Data Scale
The recurring lesson in deep learning is that large-scale high-quality data wins, but researchers are struggling to materialize this promise for agents (Xue et al., 2025). Our paper aims to solve the data problem blocking researchers from materializing this promise, and this experiment provides a valuable signal that scaling high-quality data allows small models to compete with strong LLMs from top industry labs. To proceed, we collect 20k trajectories using a Qwen 3 235B data collection policy, annotated with scores from a Qwen 3 235B judge. We then train agents based on Qwen 3 1.7B with SFT on varying scales of the data. Results in Figure 7 show that performance improves with increasing data scale, and gains scale faster on data filtered by the judge. To filter the data, we select trajectories where Judge(Success) = 1. Our top checkpoint outperforms the Qwen 3 235B data collection policy, and beats Llama 4 Maverick, a frontier LLM with 400B parameters, for which our model is 235 times smaller. The trend in the figure suggests there is room to scale further.
Figure 8: Our agents zero-shot transfer to WebVoyager. With no additional training or specialized data, our checkpoints for Qwen 3 1.7B in Section 6.1 zero-shot transfer to the WebVoyager benchmark. Trends found on our test set appear to hold for WebVoyager as well, and our top checkpoints for Qwen 3 1.7B continue to match frontier LLMs in performance for three of four judges.
6.2 Agents Transfer To New Domains
Statistically correct evaluation for deep learning models requires a test dataset that does not overlap with the training dataset, but researchers do not agree on how to implement this guideline for agents. Recent works train agents on the same websites they test on (Murty et al., 2025; Zhou et al., 2024c; Putta et al., 2024), which may obfuscate progress (Xue et al., 2025). Our next experiment shows that we can implement a stronger train-test split. Agents trained with our data can zero-shot transfer to WebVoyager (He et al., 2024) without any data from the benchmark. Results in Figure 8 show our Qwen 3 1.7B checkpoint matching strong LLMs on WebVoyager (He et al., 2024) for three of four judges, confirming trends in Section 6.1. Our ability to zero-shot transfer relatively small agents to WebVoyager (He et al., 2024) makes it likely that our pipeline leads to capable agents.
Figure 9: Our data transfers to static benchmarks. We train agents with all human data from the WebLINX and Mind2Web training sets, and resulting agents struggle to generalize to more diverse test data. Adding our data improves generalization by +149.0% for WebLINX, and +156.3% for Mind2Web.
Static Benchmarks.
To complement the online evaluation in Figure 8, we also explore how our data impacts agents trained on static benchmarks. We first train baseline agents on human demonstrations from the Mind2Web (Deng et al., 2023) and WebLINX (Lù et al., 2024) datasets. We write a preprocessor to convert our data into the expected action-observation formats these benchmarks use (which involves discarding our reasoning trace). With data converted, we compare the baseline to agents trained on a mix of 80% human data, and 20% our data, and test on (1) their official test set, and (2) 500 diverse sites from our official test set. Results in Figure 9 show that agents trained with our data perform equally well on the original test sets for these benchmark, but generalize better to our harder test set. Overall, we see +149.0% for WebLINX agents, +156.3% for Mind2Web agents, and gains in Step Accuracy on our test set are larger for the harder tasks. Additional experimental details are listed in Appendix H. On three popular benchmarks (WebVoyager, WebLINX, Mind2Web), our pipeline leads to capable agents and does not depend on human annotations.
6.3 Performance Scales With Test-time Compute
To understand the ability for web agents to scale with additional test-time compute, we ablate the number of tokens in the reasoning budget. Figure 10 shows the success rate as a function of the reasoning budget for Gemini 2.5 Flash, the top-performing agent we tested. There is a monotonic improvement in the success rate as the reasoning budget increases, and the trend suggests that performance may not be saturated with 500 reasoning tokens. Training language model agents to reason before taking actions is a promising path to better agents, and we are releasing a large reasoning dataset for multimodal agents to study this. Our dataset contains 2.2M screenshots, 2.2M reasoning traces for actions, 150k traces for judge evaluations, and led to the results in Section 6.1. The data will be linked on our website, alongside an official huggingface dataset for tasks.
Figure 10: Agents improve with a larger reasoning budget. We ablate the number of tokens in the reasoning budget for the top-performing agent, and see a monotonic improvement in the success rate as the reasoning budget increases. Gemini 2.5 Flash has a 70% success rate with a budget of 500 reasoning tokens, up from 60% for a budget of 50 tokens. The scaling of performance with the reasoning budget highlights a promising behavior in successful web agents that can be studied.
7 Conclusion
In the spirit of deep learning, we have developed an approach to efficiently harness internet data for LLM agents, and have unlocked a dynamic internet-scale environment. In building this environment, we presented a method to annotate 150k diverse sites with challenging agentic tasks, and showed how training on data from our pipeline allows small models to compete with frontier LLMs as agents, on a fraction of the budget. Our pipeline consists of a task proposer, agent, and judge driven by pretrained language models that together curate high-quality data for agents, without human intervention.
Our top checkpoint for Qwen 3 1.7B has a success rate of 56.9% on our test environment, outperforming the data collection policy Qwen 3 235B, beating the 235 times larger Llama 4 Maverick, and reaching 94.7% of the performance of Gemini 2.5 Flash, while being smaller and faster than these. Our models zero-shot transfer to WebVoyager, and scale with test-time compute. We are releasing the entire pipeline, including code, models and data, so that it may serve as a foundation for researchers to build the next generation of language model agents with internet data.
7.1 Future Work
Our work reveals several exciting directions in future work. First, our work can be scaled further: the latest Common Crawl release contains data for more than 300M sites, suggesting another 1,000 times more data could be available for agents by scaling the pipeline. In addition, we trained agents to optimize the judge scores indirectly via filtered SFT, and the judge’s high accuracy suggests that it could be optimized via reinforcement learning instead. RL is especially promising for how it can improve reasoning capabilities in agents. Finally, while the data we collect is multimodal, we focus on textual tasks in this paper, and our pipeline could be extended to produce multimodal tasks.
8 Acknowledgments
The authors recognize funding via an unrestricted gift from Amazon. In addition, the leading author was partially funded by the US Department of Defense via the NDSEG Fellowship. We thank fellow Ph.D. students who provided feedback on versions of this manuscript, including Michael Luo, Justin Wong, Tiffany Min, Murtaza Dalal, Jing-Yu Koh, Boyuan Chen, and others. We also thank faculty who provided feedback and direction for parts of this work, including Aviral Kumar, Graham Neubig, Shuyan Zhou, and others. Finally, we thank the open source community for developing systems and methods for serving LLMs efficiently, without which this paper would not be possible.
Appendix A Limitations & Safeguards
Language model agents present unique challenges and risks when applied to live tasks on the internet. For instance, agents visiting shopping sites can influence the statistics produced by analytics tools, which can impact prices on products, and product decisions from companies. Furthermore, agents seeing harmful content on the web can add that content to datasets inadvertently, and propagate harmful behaviors to future agents. We mitigate these risks in the design of the task proposal stage. We consider the risks posed to analytics tools by limiting the engagement between agents and sites. We generate only one task per website, and we limit agents to just 30 actions per site, which includes clicks, typing, dropdown selection, and more. By limiting the interaction between agents and sites, the change in website traffic generated by the InSTA pipeline is minimal (just $90$ seconds of interaction per site on average). By utilizing data from the InSTA pipeline in an offline fashion, as in Section 6 of the main paper, no additional web traffic is generated when training agents. To ensure that agents do not modify the state of the web (i.e. avoid attempting to make purchases, avoid leaving comments on posts, avoid making accounts, etc), we provide instruct the task proposer (see Figure 2) to avoid writing tasks that require the agent to interact with personal data, or user accounts.
The task proposer is instructed via the system prompt to filter out sites with harmful content, sites not intended for user access, and sites that require making an account to operate, including social media, and forum sites. There is likely a manner to safely train agents to operate user accounts, but we leave this task to future researchers. We explore the performance of the task proposer at filtering out unsuitable sites in Section 4.2, and find that all models detect unsuitable sites with a recall from $0.98$ to $1.0$, and accuracy up to 97%, suggesting our filter is reliable. Sites used to benchmark the performance of the safety filter are discussed in Appendix E, and thoroughly test the safety filter.
To remove Personally Identifiable Information (PII) from the data used for training agents, we include scrubadub, an industry standard PII removal tool for python developed by Leap Beyond, a data consultancy based in the European Union. Our pipeline has an argument that toggles the usage of scrubadub to remove PII from all website data, and we recommend this option be set.
Appendix B Ethical Considerations
One important ethical consideration when harnessing internet data is to carefully handle copyrighted, private, and sensitive materials. The internet contains vast amounts of PII, which should be avoided when training models. We address this in two ways. First, we instruct the task proposer to filter out sites that may contain PII, including social media websites, and forums. Second, our pipeline has an argument that toggles the usage of scrubadub to remove PII, and we recommend this option be set. To prevent copyrighted materials from being used for training, we do not allow the agent to copy any books, documents, videos, audio, or files. The information present in the HTML DOM is publically available, and is converted into a sparse Markdown summary that focuses on interactive elements, and significantly transforms the underlying webpage in a manner that is not reversible.
These steps significantly reduce, but do not completely eliminate the risk that private, and sensitive materials are shown to the agent, and methods for detecting, replacing, and removing such materials from text and images remains an important task for researchers working on safety.
Appendix C Broader Impacts
As their capabilities broaden, language models are being used to interface with real-world systems. This shift comes with several benefits and risks. Agents that operate your computer to aid in work tasks can significantly boost productivity for certain workers, but can displace others whose jobs have been fully automated. Agents that operate browsers to complete personal tasks provide convenience, but expose a new attack vector where rogue agents perform unintended actions. Certain risks can be mitigated with proper safeguards, such as post-processing data to prevent jail-breaking, but other risks are existential, harder to address purely with new agent research, and may require policy changes.
Appendix D Agents.txt & Standards For Internet Agents
Akin to robots.txt directives, website creators should have a standard format to specify how internet agents are allowed to interact with their websites, and what information on webpages agents are allowed to see. Desireable controls include rate limits for interactions, limits for maximum numbers of interactions, restrictions to allow agents to interact with certain pages and not others, and restrictions on the kind of data on webpages that agents are allowed to observe (achieved via tagging elements to hide their content from agents). In addition to restricting the data available to agents, website creators should have the ability to specify locations for “playgrounds” that replicate certain key functions of their site with virtual tasks and simulated data that are intended to teach agents how to operate their site while directing traffic from agents away from user-facing pages.
Appendix E More Details On Task Generation
You are helping us create tasks for a web navigation system. We will tell you the domain of a website. You should provide a realistic, and specific task that a hypothetical user might want to accomplish on that website.
## Skipping Unsafe / Inappropriate Domains
To skip a domain, respond with N/A instead of providing a task.
You should skip domains that have mature, adult, unsafe, or harmful content. If you are unsure whether a domain is safe, please skip it. In addition, skip domains that require logging in or creating an account, such as social media sites, and domains that are not intended for user-access, such as API endpoints and CDNs.
## Here are some domains to provide tasks for:
* www.amazon.com‘: Find the price of the 24in LG Ultragear Monitor.‘
* www.wikipedia.org‘: Look up the history of the Eiffel Tower on Wikipedia.‘
## Here are some domains to skip:
* fbcdn.net‘: N/A
* api.github.com‘: N/A
Tasks should not require external knowledge, not modify the state of the web, and should not require logging in or creating an account. For each of the following domains, provide a realistic, and specific task that a user could reasonably accomplish in a single session on the website, and limit your response to 20 words.
Figure 11: System prompt for the exploration phase of task generation. We design the system prompt for task generation to detect and remove unsafe websites. This prompt ensures that tasks are passive, and do not modify content on a website. Refer to the next figures for the in-context examples used for the task proposer, and the system prompt used in the feedback step.
We provide the system prompt used in the first phase of the task generation loop in Figure 11. This prompt was provided to Llama 3.1 70B, GPT-4o, and Gemini 1.5 Pro to generate tasks and filter sites unsuitable for annotation in Section 4. We carefully designed this system prompt to enforce that generated tasks are passive, and do not modify content on a website. In addition to this system prompt, we employed a list of 100 hand-picked in-context examples of website URLs and appropriate tasks, which are provided in the following JSON list. When querying an LLM, we randomly sample 16 in-context examples from the list, and provide only these examples to the LLM to generate a task to guide exploration of the site. This improves diversity in the exploration phase.
"domain": "archive.org",
"task": "Identify the oldest book available in the public domain on this site."
"domain": "arxiv.org",
"task": "Retrieve the latest preprint paper on machine learning."
"domain": "wikibooks.org",
"task": "Find a freely available textbook on linear algebra."
"domain": "wiktionary.org",
"task": "Get the definition and etymology of the word serendipity’."
"domain": "openlibrary.org",
"task": "Locate an ebook about classic literature that is available for borrowing."
"domain": "openculture.com",
"task": "Find a free online course on ancient history."
"domain": "theguardian.com",
"task": "Retrieve an article discussing recent trends in renewable energy."
"domain": "medium.com",
"task": "Identify a highly rated blog post on productivity hacks."
"domain": "goodreads.com",
"task": "Find the most popular book related to neuroscience."
"domain": "wired.com",
"task": "Retrieve an article about the latest advancements in wearable technology."
"domain": "data.gov",
"task": "Identify the latest government dataset on climate change."
"domain": "kaggle.com",
"task": "Find a well-documented data science competition on image recognition."
"domain": "gov.uk",
"task": "Locate the latest UK government report on healthcare."
"domain": "unsplash.com",
"task": "Find a high-resolution image of the Milky Way Galaxy."
"domain": "pexels.com",
"task": "Retrieve a popular photo tagged with nature’."
"domain": "creativecommons.org",
"task": "Find an article explaining Creative Commons licensing types."
"domain": "pypi.org",
"task": "Retrieve the most downloaded Python package for data analysis."
"domain": "huggingface.co",
"task": "Identify a popular machine learning model on this platform."
"domain": "sciencenews.org",
"task": "Find the most recent article on the health impacts of air pollution."
"domain": "mit.edu",
"task": "Retrieve a publicly available research paper on quantum computing."
"domain": "springer.com",
"task": "Identify the latest edition of a Springer book on robotics."
"domain": "jstor.org",
"task": "Find a research paper discussing the history of the Internet."
"domain": "biorxiv.org",
"task": "Retrieve the most recent bioRxiv preprint on CRISPR technology."
"domain": "medrxiv.org",
"task": "Find a public health preprint related to COVID-19."
"domain": "commons.wikimedia.org",
"task": "Retrieve a high-resolution image of the Eiffel Tower."
"domain": "scholar.google.com",
"task": "Find the most cited article by a specific researcher."
"domain": "plos.org",
"task": "Locate the latest research paper on gene editing published here."
"domain": "flickr.com",
"task": "Find a photo that has been released under a Creative Commons license."
"domain": "datacite.org",
"task": "Retrieve metadata for a dataset related to environmental studies."
"domain": "orcid.org",
"task": "Find the ORCID ID of a well-known researcher in AI."
"domain": "zotero.org",
"task": "Retrieve an article discussing citation management tools."
"domain": "github.com",
"task": "Find the most starred repository on deep learning."
"domain": "figshare.com",
"task": "Retrieve an open dataset on climate patterns."
"domain": "zenodo.org",
"task": "Find the latest publication on open science practices."
"domain": "worldcat.org",
"task": "Locate a catalog entry for a rare book on botany."
"domain": "biodiversitylibrary.org",
"task": "Retrieve a scanned copy of an 18th-century botanical illustration."
"domain": "genome.gov",
"task": "Find the latest update on the Human Genome Project."
"domain": "merriam-webster.com",
"task": "Retrieve the definition and usage of the word quantum’."
"domain": "stanford.edu",
"task": "Find the most recent online lecture on artificial intelligence."
"domain": "edx.org",
"task": "Retrieve a TED Talk on leadership in technology."
"domain": "ted.com",
"task": "Find the latest ocean temperature data available."
"domain": "noaa.gov",
"task": "Retrieve a dataset related to consumer behavior."
"domain": "data.world",
"task": "Find a course on data visualization."
"domain": "curious.com",
"task": "Retrieve a well-cited article on the psychological impact of social media."
"domain": "theconversation.com",
"task": "Identify a recent research paper on biodiversity conservation."
"domain": "nature.com",
"task": "Retrieve the latest article on genomics research."
"domain": "pnas.org",
"task": "Find a science news article on robotics advancements."
"domain": "sciencedaily.com",
"task": "Identify the top story on global health issues."
"domain": "bbc.com",
"task": "Retrieve a recent podcast episode about space exploration."
"domain": "npr.org",
"task": "Locate the most recent update on the global biodiversity status."
We also provide the system prompt used in the second phase of the task generation loop, where trajectories from agents are fed back to the task proposer, which generates a harder, grounded task. This prompt instructs the task proposer to create a challenging task based on how an expert user could be expected to use the shown website. The task proposer also predicts a list on intermediate steps that can be used as a hint for agents, and a success criteria that can be used to improve the verifier.
You are a helpful assistant designing tasks for a web automation script. I will show you previous runs of the script, including previous tasks, webpages, actions, and performance reviews, formatted in markdown. Help me design *challenging* new tasks.
## Formatting The Proposed Task
Format your task in the following JSON schema:
‘‘‘json
"proposed_task": str,
"steps": List[str],
"criteria": str
Here is what each key means:
- proposed_task‘: A specific, challenging task that an expert user might leverage this website to complete.
- Must not require making an account, logging in, submitting personal information, making a purchase, or placing an order.
- steps‘: Steps an expert user would follow to complete the proposed task.
- criteria‘: The required answer, and criteria to determine if the task was completed.
## Example Tasks For Inspiration
Suppose you want to design a task around the C-to-C Hose-Shut-Off Valve on awg-fittings.com’:
‘‘‘json
"proposed_task": "What is the C-to-C Hose-Shut-Off Valve length in mm?",
"steps": [
"Navigate to awg-fittings.com’",
"Open the product catelog for fittings",
"Locate the product listing for the C-to-C Hose-Shut-Off Valve",
"Find the product length in mm, and respond with that length in the answer"
"criteria": "The answer should include the specific length of ’237 mm for this product"
Suppose you want to design a task around the document The Angora cat; how to breed train and keep it on biodiversitylibrary.org’:
‘‘‘json
"proposed_task": "Open a scanned copy of The Angora cat; how to breed train and keep it’.",
"steps": [
"Navigate to biodiversitylibrary.org’",
"Search for The Angora cat; how to breed train and keep it in the search bar",
"Click on the title of the document in the search results",
"Confirm the correct document is displayed in an embedded PDF reader"
"criteria": "The final webpage should display the correct document in an embedded PDF reader"
Suppose you want to design a task around the Generative Adversarial Networks paper on scholar.google.com’:
‘‘‘json
"proposed_task": "How many citations does the paper Generative Adversarial Networks have?",
"steps": [
"Navigate to scholar.google.com’",
"Search for Generative Adversarial Networks in the search bar",
"Locate the correct paper in the search results",
"Find an up-to-date citation count, and respond with that count in the answer"
"criteria": "The answer should include an up-to-date citation count, which is ’80613’ as of April 2025"
Suppose you want to design a task around the word serendipity on wiktionary.org’:
‘‘‘json
"proposed_task": "What is the definition and etymology of the word serendipity’?",
"steps": [
"Navigate to wiktionary.org’",
"Search for serendipity in the search bar",
"Find the definition and etymology sections of the serendipity page",
"Summarize the contents of these sections in the answer"
"criteria": "The answer should mention Serendip (or Serendib), coined by English writer and politician Horace Walpole in 1754"
Thanks for helping me design challenging new tasks, please follow the instructions carefully. Start your response with an analysis for how an expert user would leverage this website, followed by a step-by-step breakdown of your proposed task, and finally, enter your task in the JSON format. Respond in 500 words.
E.1 Details For Safety Experiment
Using these prompts for task generation, we remove unsafe websites. To evaluate the performance of our filter, we employed a set of 100 curated websites, where 50 are manually verified as safe, and 50 are manually verified as unsafe based on the filtering conditions. These sites were chosen to span popular sites that typical users are likely familiar with, and less popular websites that may be underrepresented in LLM training data.
safe_sites_list = [’dhss.mo.gov’, dizionari.corriere.it’, southgippsland.vic.gov.au’, ds.iris.edu’, lobbycontrol.de’, ’4rsmokehouse.com’, barnsleyfc.co.uk’, wiwi.uni-wuerzburg.de’, uplandca.gov’, lsus.edu’, wpcode.com’, webopedia.internet.com’, tamko.com’, premierchristian.news’, genome.jgi.doe.gov’, burgerking.ca’, thehugoawards.org’, radio.fm’, thevinyldistrict.com’, unilang.org’, raywhitegroup.com’, grapevinetexas.gov’, sanfrancisco.cbslocal.com’, hyde-design.co.uk’, breastcancerfoundation.org.nz’, ludwigsburg.de’, ignitionrobotics.org’, deliverit.com.au’, kodokan.org’, clickstay.com’, searchdatamanagement.techtarget.com’, oceanario.pt’, wentworthpuzzles.com’, catholicworldreport.com’, quizlet.com’, innovation.nhs.uk’, synonyms.reverso.net’, news.siemens.co.uk’, readability-score.com’, co.modoc.ca.us’, cityofmyrtlebeach.com’, loire.gouv.fr’, lawphil.net’, saem.org’, parmigianoreggiano.it’, engaging-data.com’, itf-tkd.org’, aka.education.gov.uk’, ub.uni-kl.de’, mottchildren.org’]
unsafe_sites_list = [’refer to the code release’]
We provide the list of unsafe sites in the code to keep this manuscript clean. Using these lists, we conduct an experiment measuring the accuracy, precision, and recall of our safety filter for detecting sites that are not suitable for training agents. In this experiment, we run the initial exploration phase of the task proposer with the same prompts used in the main experiments (shown in the last section), and we consider a website to be marked positive for unsafe content if the task proposer generates “N/A” rather than a task. We then calculate accuracy, precision, and recall for various LLMs.
E.2 Details For Reliability Experiments
We evaluated the verifiable rate of tasks generated by the initial phase of task generation manually. For a set of 100 randomly sampled websites marked as safe by the task proposer in the initial phase, we attempted to complete the generated task, and checked that enough information is present on the website that a solution can be verified. In total, we annotated 300 tasks for Table 2 in 6 hours, and provide the 100 website URLs used in this experiment in the following code block.
reliability_sites_list = [’godaddy.com’, chrome.google.com’, apple.com’, support.cloudflare.com’, support.apple.com’, edition.cnn.com’, go.microsoft.com’, google.de’, w3.org’, yandex.ru’, bfdi.bund.de’, microsoft.com’, apps.apple.com’, networksolutions.com’, support.mozilla.org’, yelp.com’, cnn.com’, ec.europa.eu’, developer.mozilla.org’, icann.org’, books.google.com’, globenewswire.com’, onlinelibrary.wiley.com’, gnu.org’, slideshare.net’, metacpan.org’, porkbun.com’, oag.ca.gov’, spiegel.de’, linuxfoundation.org’, help.opera.com’, mayoclinic.org’, podcasts.apple.com’, nhs.uk’, addons.mozilla.org’, google.fr’, pewresearch.org’, finance.yahoo.com’, weforum.org’, g2.com’, savethechildren.org’, news.com.au’, biblia.com’, yr.no’, engadget.com’, microsoftstore.com’, ema.europa.eu’, theintercept.com’, princeton.edu’, foodandwine.com’, sfgate.com’, voguebusiness.com’, ourworldindata.org’, livingwage.org.uk’, cms.law’, msdmanuals.com’, websitesetup.org’, support.xbox.com’, treehugger.com’, tripadvisor.com.pe’, mondragon.edu’, greenparty.ca’, aaojournal.org’, restaurantpassion.com’, iwillteachyoutoberich.com’, moneyconvert.net’, gesundheitsinformation.de’, ovc.uoguelph.ca’, zdnet.be’, oxfordamerican.org’, snackandbakery.com’, journals.uic.edu’, confused.com’, standards.globalspec.com’, onlyinyourstate.com’, ahsgardening.org’, wyze.com’, nornickel.ru’, viessmann.fr’, benetton.com’, firecomm.gov.mb.ca’, executedtoday.com’, eukn.eu’, fraeylemaborg.nl’, verizon.com/about/news-center’, orthodoxalbania.org’, cheapjoes.com’, bake-eat-repeat.com’, plattformpatientensicherheit.at’, hifinews.com’, cellsignal.com’, thenotariessociety.org.uk’, chosenfoods.com’, westerndressageassociation.org’, pridesource.com’, northtacomapediatricdental.com’, strade-bianche.it’, pvdairport.com’, institute.sandiegozoo.org’, raintaxi.com’]
E.3 Automatic Task Categorization
We employ Llama 3.1 70B to categorize tasks. We prompt Llama 3.1 70B with the system prompt in Figure 12 to assign a category in 3 words or less to encourage simple categories. Categories have 16.9 tasks on average, and 953 categories have more than the mean, while 7741 have less than the mean. There is occasional overlap between categories, which can be observed in Figure 13, but for the purposes of understanding performance by category, overlap is acceptable provided categories have sufficiently large numbers of tasks, and performance per category can be accurately calculated. We provide our task categorization script in the official code release.
You are a helpful scientific assistant categorizing tasks on the web. You will observe a domain and web navigation task, and you should provide a concise categorization of the task in 3 words or less. For example, if the domain is "google.com" and the task is "find a recipe for mashed potato", you may categorize the task as "recipe search".
## Task Format
Here is the format for the task:
[domain]: [task]
Here is what each part means:
‘[domain]‘: The domain of the website you are observing.
‘[task]‘: The task a user is trying to accomplish on the website.
## Response Format
Respond with a category name for the task in 3 words or less, and provide only the category name, do not provide an explanation or justification for the categorization.
Here is the next task, please follow the instructions carefully.
Figure 12: System prompt for task categorization. We employ Llama 3.1 70B to automatically label task categories for our dataset. We prompt the LLM to assign categories in 3 words or less, and set the sampling temperature to $0.5$ to encourage predictions to use more consistent language. Using these categories, we seek to understand agent performance by category.
Appendix F Understanding Agent Capabilities & Limitations
Figure 13: Largest categories for task generation. We categorize 150k tasks generated by our pipeline in Section 4, and visualize the number of tasks in the largest 70 categories. Top categories include article search, news search, recipe search, and product lookup. The top 12 task categories have more than 1600 tasks assigned to each of them, the mean number of tasks per category is 16.9, and 89% of categories (7741 in total) have fewer than the mean number of tasks.
Figure 14: Most solved categories for task generation. We explore the completion rates for the top categories of tasks generated by our pipeline. We restrict our focus to categories where at least 100 tasks are assigned, and plots the success rates for the top 70 categories. Results show that 22 of these categories are solved with more than a 50% rate with zero-shot agents based on Llama 3.1 70B.
To complement the analyses presented in Section 5, we explore the categories of tasks that agents succeed at most frequently. Shown in Figure 14, we plot the average judge success probability prediction $\mathbf{r}_{T}$ versus task category for the top 70 most successful categories that have at least 100 tasks assigned to them. Based on the figure, top categories include searching for contact information, finding hours of operation, looking up biographical information, obtaining current weather forecasts, and conducting health research. Based on these results, the top 22 categories are solved with an average success probability $>0.5$ using zero-shot agents based on Llama 3.1 70B. As stronger models are developed, the success rates for agents running in our pipeline are likely to improve, and the quality of the data we generate will jointly improve.
In addition to studying the best-performing categories, we also explore the limitations of current agents via their least successful categories. Shown in Figure 15, we select the bottom 70 categories via their average judge success probability for categories with at least 100 tasks assigned. Many of these categories require agents to remember and reason about previous interactions, such as the product comparison category. For this category, an agent must review several products, and compare details from memory. In these cases, access to a note-taking tool may improve performance. Additionally, certain task categories involve requests that are not feasible given the limitations of the Playwright API, including categories for downloading reports / manuals, and opening files. While these tasks are not currently feasible, providing agents with a fully-operable virtual computer environment with applications pre-installed could unlock these abilities in future work.
Figure 15: Least successful categories for internet-scale task generation. Similar to the previous figure, we explore the rates of task completion for the bottom 70 categories that have at least 100 tasks assigned to them. While the majority of the least successful categories have success rates greater than 20%, performance drops as low as 5%. Many of the categories shown in the plot above involve actions that are not feasible given the current limitations of the Playwright API, and may be possible in future work that extends agents to a fully-operable virtual computer environment. In addition, better LLM backbones are likely to improve performance.
Appendix G Agent & Judge System Prompts
We provide the system prompt that powers the agent in this paper. This prompt is released in the code, alongside a fast HTML to Markdown processor we built. The agent prompt is carefully designed to elicit reasoning capabilities, and the experiment in Figure 10 shows the prompt is successful.
You are helping me complete tasks by operating a web browser. I will share the current task, and a sequence of webpages and actions from previous steps.
## Action Instructions
Based on the information we discovered so far, and the progress we made in previous steps, you are helping me determine the next action.
You will provide an action as JSON in a fenced code block:
‘‘‘json
"action_key": str,
"action_kwargs": dict,
"target_element_id": int | null
Actions have the following components:
- action_key‘: The name of the selected action.
- action_kwargs‘: A dictionary of arguments for the action.
- target_element_id‘: An optional id for the element to call the action on.
## Action Definitions
Ive prepared an API documentation below that defines the actions we can use to complete the task.
### Click Action Definition
- click‘: Click on an element specified by target_element_id‘.
### Example Click Action
Suppose you want to click ‘[id: 5] Sales link‘:
‘‘‘json
"action_key": "click",
"action_kwargs": {},
"target_element_id": 5
### Hover Action Definition
- hover‘: Hover over an element specified by target_element_id
### Example Hover Action
Suppose you want to hover over ‘[id: 2] Company Logo image‘:
‘‘‘json
"action_key": "hover",
"action_kwargs": {},
"target_element_id": 2
### Scroll Action Definition
- scroll‘: Scroll the page by delta_x pixels to the right and delta_y pixels down.
- delta_x‘: The number of pixels to scroll to the right.
- delta_y‘: The number of pixels to scroll down.
### Example Scroll Action
Suppose you want to scroll down the page by 300 pixels:
‘‘‘json
"action_key": "scroll",
"action_kwargs": {
"delta_x": 0,
"delta_y": 300
"target_element_id": null
### Fill Action Definition
- fill‘: Fill an input element specified by target_element_id with text.
- value‘: The text value to fill into the element.
### Example Fill Action (Text Input)
Suppose you want to fill ‘[id: 13] "Name..." (Enter your name text input)‘ with the text John Doe‘:
‘‘‘json
"action_key": "fill",
"action_kwargs": {
"value": "John Doe"
"target_element_id": 13
### Example Fill Action (Range Slider)
Suppose you want to set ‘[id: 71] "\$250 (5)" (range slider min: 0 max: 50 step: 1)‘ to the value of \$1000‘. The slider has a range of 0 to 50 with a step of 1, and the value is currently set to ‘5‘. You must translate the desired \$1000 to the correct underlying value of ‘20‘:
‘‘‘json
"action_key": "fill",
"action_kwargs": {
"value": "20"
"target_element_id": 71
### Select Action Definition
- select‘: Select from a dropdown element specified by target_element_id‘.
- label‘: The option name to select in the element.
### Example Select Action
Suppose you want to select the option red from ‘[id: 67] "blue" (color select from: red, blue, green)‘:
‘‘‘json
"action_key": "select_option",
"action_kwargs": {
"label": "red"
"target_element_id": 67
### Set Checked Action Definition
- set_checked‘: Check or uncheck a checkbox specified by target_element_id‘.
- checked‘: Boolean value to check or uncheck the checkbox.
### Example Set Checked Action
Suppose you want to check ‘[id: 21] "I agree to the terms and conditions" (checkbox)‘:
‘‘‘json
"action_key": "set_checked",
"action_kwargs": {
"checked": true
"target_element_id": 21
### Go Back Action Definition
- go_back‘: Go back to the previous page (‘target_element_id must be null).
### Example Go Back Action
Suppose you want to go back to the previous page:
‘‘‘json
"action_key": "go_back",
"action_kwargs": {},
"target_element_id": null
### Goto Action Definition
- goto‘: Navigate to a new page (‘target_element_id must be null).
- url‘: The URL of the page to navigate to.
### Example Goto Action
Suppose you want to open the DuckDuckGo search engine:
‘‘‘json
"action_key": "goto",
"action_kwargs": {
"url": "https://www.duckduckgo.com"
"target_element_id": null
### Stop Action Definition
- stop‘: Stop when the task is complete, and report your progress.
- answer‘: Optional answer sent back to me.
### Example Stop Action
Suppose the task is complete, and you want to stop and report your progress:
‘‘‘json
"action_key": "stop",
"action_kwargs": {
"answer": "The desired task is now complete."
"target_element_id": null
## Formatting Your Response
Write a 200 word revised plan based on new information we discovered, and progress we made in previous steps. After your response, provide the next action as JSON in a fenced code block.
We also provide the system prompt used by the judge. The system prompt instructs the judge to predict JSON within a fenced code block that contains a “success” key, an “efficiency” key, and a “self_correction” key. The success key represents a score from 0 to 1 that estimates the probability the task is successfully completed. The efficiency key represents a score from 0 to 1 that estimates the probability the agent has taken the most efficient path to solve the task. The self correction key represents a score from 0 to 1 that estimates the probability that the agent has demonstrated self corrective behaviors during its completion of the task. These behaviors include when the agent backtracks to a more promising state, re-plans when new information is discovered relevant to the task, and recognizes its own mistakes. These are generally behaviors we expect from successful agents, but for this paper we only filter by the success key to select training data for agents.
You are helping me evaluate a browser automation script. I will share a task provided to the script, and a sequence of webpages and actions produced by the script.
## The Action Format
The script produces actions as JSON in a fenced code block:
‘‘‘json
"action_key": str,
"action_kwargs": dict,
"target_element_id": int
Actions have the following components:
- action_key‘: The name of the selected action.
- action_kwargs‘: A dictionary of arguments for the action.
- target_element_id‘: An optional id for the element to call the action on.
## Action Definitions
Ive prepared an API documentation below that defines the actions the script can use to complete the task.
### Click Action Definition
- click‘: Click on an element specified by target_element_id‘.
### Example Click Action
Here is an example where the script clicks ‘[id: 5] Sales link‘:
‘‘‘json
"action_key": "click",
"action_kwargs": {},
"target_element_id": 5
### Hover Action Definition
- hover‘: Hover over an element specified by target_element_id
### Example Hover Action
Here is an example where the script hovers over ‘[id: 2] Company Logo image‘:
‘‘‘json
"action_key": "hover",
"action_kwargs": {},
"target_element_id": 2
### Scroll Action Definition
- scroll‘: Scroll the page by delta_x pixels to the right and delta_y pixels down.
- delta_x‘: The number of pixels to scroll to the right.
- delta_y‘: The number of pixels to scroll down.
### Example Scroll Action
Here is an example where the script scrolls down the page by 300 pixels:
‘‘‘json
"action_key": "scroll",
"action_kwargs": {
"delta_x": 0,
"delta_y": 300
"target_element_id": null
### Fill Action Definition
- fill‘: Fill an input element specified by target_element_id with text.
- value‘: The text value to fill into the element.
### Example Fill Action (Text Input)
Here is an example where the script fills ‘[id: 13] "Name..." (Enter your name text input)‘ with the text John Doe‘:
‘‘‘json
"action_key": "fill",
"action_kwargs": {
"value": "John Doe"
"target_element_id": 13
### Example Fill Action (Range Slider)
Here is an example where the script sets ‘[id: 71] "\$250 (5)" (range slider min: 0 max: 50 step: 1)‘ to the value of \$1000‘. This slider has a range of 0 to 50 with a step of 1, and the value is currently set to ‘5‘. The script translates the desired \$1000 to the correct underlying value of ‘20‘:
‘‘‘json
"action_key": "fill",
"action_kwargs": {
"value": "20"
"target_element_id": 71
### Select Action Definition
- select‘: Select from a dropdown element specified by target_element_id‘.
- label‘: The option name to select in the element.
### Example Select Action
Here is an example where the script selects the option red from ‘[id: 67] "blue" (color select from: red, blue, green)‘:
‘‘‘json
"action_key": "select_option",
"action_kwargs": {
"label": "red"
"target_element_id": 67
### Set Checked Action Definition
- set_checked‘: Check or uncheck a checkbox specified by target_element_id‘.
- checked‘: Boolean value to check or uncheck the checkbox.
### Example Set Checked Action
Here is an example where the script checks ‘[id: 21] "I agree to the terms and conditions" (checkbox)‘:
‘‘‘json
"action_key": "set_checked",
"action_kwargs": {
"checked": true
"target_element_id": 21
### Go Back Action Definition
- go_back‘: Go back to the previous page (‘target_element_id must be null).
### Example Go Back Action
Here is an example where the script goes back to the previous page:
‘‘‘json
"action_key": "go_back",
"action_kwargs": {},
"target_element_id": null
### Goto Action Definition
- goto‘: Navigate to a new page (‘target_element_id must be null).
- url‘: The URL of the page to navigate to.
### Example Goto Action
Here is an example where the script opens DuckDuckGo search:
‘‘‘json
"action_key": "goto",
"action_kwargs": {
"url": "https://www.duckduckgo.com"
"target_element_id": null
### Stop Action Definition
- stop‘: Stop when the task is complete, and report the progress.
- answer‘: Optional answer from the script.
### Example Stop Action
Here is an example where the script stops and reports its progress:
‘‘‘json
"action_key": "stop",
"action_kwargs": {
"answer": "The desired task is now complete."
"target_element_id": null
## Evaluation Instructions
Based on the progress of the script, you are helping me determine if the desired task has been completed successfully.
You will provide scores as JSON in a fenced code block:
‘‘‘json
"success": float,
"efficiency": float,
"self_correction": float
### Score Definitions
- success‘: Your confidence the desired task has been completed successfully.
- range: 0.0 (not possible) to 1.0 (absolutely certain).
- efficiency‘: Your confidence the script has taken the most efficient path to complete the task.
- range: 0.0 (not possible) to 1.0 (absolutely certain).
- self_correction‘: Your confidence the script has demonstrated self-corrective behaviors during its completion of the task. These behaviors include backtracking to a more promising state, replanning when new information is discovered, and recognizing its own mistakes.
- range: 0.0 (not possible) to 1.0 (absolutely certain).
Write a 300 word analysis that establishes specific criteria to rigorously evaluate whether the task was completed, followed by which criteria the script has satisfied. After your response, provide your scores as JSON in a fenced code block.
Appendix H Details For Training Agents
To understand the utility of data we obtained, we train agents and test on four relevant benchmarks: InSTA, WebVoyager (He et al., 2024), Mind2Web (Deng et al., 2023), WebLINX (Lù et al., 2024). In particular, our test set consists of a held-out set of 3,000 websites and tasks produced by the task generation feedback loop. Note these websites are not present in the training set. For WebVoyager (He et al., 2024), we transfer agents trained on our data zero-shot to 643 tasks on 15 websites WebVoyager (He et al., 2024). The websites in the WebVoyager benchmark are not present in the 20k trajectories we collected in Section 6.1 we used for training agents. For this experiment, we fine-tuned models based on Qwen 3 1.7B with a maximum sequence length of 16,384 tokens, and the most recent 5 observations, and actions in the context. We employed full fine-tuning on this model, with Adam, learning rate of 5e-5, batch size of 32, bfloat16, and other parameters kept as the PyTorch defaults for Adam. Each model was trained using one epoch, a linear warm-up corresponding to the first 0.01 steps of training, and a linear decay to 6e-5 afterwards.
To filter data, we select trajectories that were scores as Judge(Success) = 1, which corresponded to 10.5k of 20k trajectories produced by the Qwen 3 235B data collection policy. Scores for filtering were produced by a Qwen 3 235B judge. We employed a simple filtering strategy that only considers the success score from the judge, and no other filtering conditions were used. Note the judge also predicts efficiency, and self correction scores, which could likely also help select the best data for training, but we did not explore filtering by these scores in this work.
For experiments on static benchmarks, we fine-tune google/flan-t5-large for Mind2Web, and meta-llama/Llama-3.1-8B-Instruct for WebLINX using official fine-tuning code released with corresponding benchmarks. We employ identical training hyperparameters to those used by Lù et al. (2024) for Llama in their official training code and Deng et al. (2023) for Flan to ensure that our results are comparable to previous work. Section 6.2 reports performance on the official test_web split of the WebLINX benchmark, and the official test_website split of the Mind2Web benchmark, where agents are tested on previously unobserved websites. The websites in these static benchmarks were not present in the dataset produced by InSTA for this experiment, to ensure fairness.
Appendix I Hyperparameters
We provide a list of the hyperparameters used in this work in Table 3. Hyperparameters for mixing our data with human data on static benchmarks are selected to mirror prior work in synthetic data (Trabucco et al., 2024), and to adhere to standard hyperparameters for WebLINX (Lù et al., 2024), and Mind2Web (Deng et al., 2023). We train using all available human data on these benchmarks, and add trajectories filtered using the previously discussed methodology, sampled at a 20% rate in the data-loader compared to an 80% rate for human data.
Hyperparameter Name Value
Models Used For Agents Qwen/Qwen3-1.7B
Qwen/Qwen3-235B-A22B
meta-llama/Llama-4-Maverick-17B-128E-Instruct
meta-llama/Llama-3.1-70B-Instruct
meta-llama/Llama-3.3-70B-Instruct
google/gemini-2.5-flash
Models Used For Judges Qwen/Qwen3-235B-A22B
meta-llama/Llama-4-Maverick-17B-128E-Instruct
meta-llama/Llama-3.1-70B-Instruct
meta-llama/Llama-3.3-70B-Instruct
google/gemini-2.5-flash
google/gemini-1.5-pro
openai/gpt-4.1-mini
openai/gpt-4o
Common Crawl PageRank cc-main-2024-apr-may-jun-host-ranks.txt.gz
Number of sites before filtering $1,000,000$
Number of tasks after filtering $146,746$
Max Tokens Per Observation $2,048$
Max Tokens Per Agent Trace $1,024$
Max Tokens Per Judge Trace $1,024$
Max Tokens Per Task Proposer Trace $1,024$
Last Steps Per Agent Context 5
Last Steps Per Judge Context 5
Last Steps Per Task Proposer Context 5
Task Proposer Feedback Loops 1
LLM Sampling Temperature 0.5
LLM Sampling Top P 1.0
LLM Sampling Top K default
Fine-tuned LLM in Section 6.1 Qwen/Qwen3-1.7B
InSTA Training Epochs 1
InSTA Batch Size $32$
InSTA Learning Rate 5e-5
InSTA Optimizer Adam
Mind2Web LLM google/flan-t5-large
Mind2Web Training Iterations $11,505$
Mind2Web Batch Size $32$
Mind2Web Learning Rate 5e-5
Mind2Web Optimizer Adam
WebLINX LLM meta-llama/Llama-3.1-8B-Instruct
WebLINX Training Iterations $10,000$
WebLINX Batch Size $16$
WebLINX Learning Rate 5e-5
WebLINX Optimizer Adam
Data Filtering Condition Judge(Success) = 1
Human Data Sampling Probability $p_{\text{real}}$ 80%
Table 3: Hyperparameters used in our paper. We organize hyperparameters into seven sections, for the names of LLMs used as agents in the paper, the names of LLMs used as judges in the paper, the hyperparameters used for data collection, the sampling parameters for LLMs, the training parameters for static benchmarks, and the filtering and data mixing hyperparameters.
Appendix J Cost Analysis For Llama 3.1 70B
To understand the significant reduction in cost that we obtain by running LLMs locally to generate data, we analyze the number of tokens processed by the LLM, and compute an expected cost if this were served using proprietary models. As the analysis shows, using Llama 3.1 70B is a feasible option for running agents at this large scale, and results in the paper show that this choice of LLM backbone does not compromise performance. We have deep gratitude to the Llama team at Meta, and the Qwen team at Alibaba for working to make developments in language modeling available to the research community at no cost. We see up to a 95% reduction in cost with these models.
Variable Name Value
Number of tasks $146,746$
Average tokens per observation $1,024$
Max observations per agent context window $5$
Average agent / judge response size $512$
Max tokens per system prompt $1,024$
Average steps per task $15$
Estimated tokens processed by the agent 14.65B tokens
Tokens processed by the judge 1.35B tokens
Total tokens processed 16.00B tokens
Expected API cost for GPT-4.1 \$ 32,000.00
Expected API cost for Gemini 2.5 Pro \$ 20,000.00
Expected AWS compute cost for serving Llama 3.1 70B
       (3,840 v100 GPU hours using spot instances) \$ $1,575.70$
Percent saved using Llama 3.1 70B $[95.08,92.12]$
Table 4: Cost analysis for different LLM models in the fully-scaled pipeline. This table provides statistics for the number of tokens that were processed by our pipeline, and why serving using a local LLM engine like vLLM is important for bringing down costs.